# An Order of Magnitude Time Complexity Reduction for Gaussian Graphical Model Posterior Sampling Using a Reverse Telescoping Block Decomposition

### v1

@@ 4 Numerical Experiments
We compare the proposed reverse telescoping (RT) MCMC sampler with the cyclical MCMC sampler under the graphical horseshoe (GHS, Li et al.,, 2019) and graphical horseshoe-like (GHSL, Sagar et al., 2024a, ) priors in terms of estimation results and computational efficiency. To ensure $n=o(p^{1/2})$, we set $n\approx 24\log p$ for $p=100,150,200,250,300$; resulting in the following problem dimensions: $(n,p)\in\{(110,100),(120,150),(127,200),(132,250),(137,300)\}$. Recall the theory suggests in this regime, the RT sampler achieves $\mathcal{O}(\max\{n^{2}p^{2},p^{3}\})=\mathcal{O}(p^{3})$ per-iteration complexity, whereas the complexity of the cyclical sampler remains $\mathcal{O}(p^{4})$.
@@
The true precision matrix $\boldsymbol{\Theta}_{0}$ is set as in Li et al., (2019) and Sagar et al., 2024b . The diagonal elements are set to one and the off-diagonal elements are set to follow the cliques positive structure, which means the $p$ rows and columns are partitioned into $K$ disjoint groups $\{\mathbf{G}_{k}\}_{k=1}^{K}$ and we select 3 members within each group, $g_{k}\subset\mathbf{G}_{k}$. the off-diagonal entry is set to 0.45-0.45 (corresponding to positive partial correlations) if $i\neq j$ and $i,j\in g_{k}$, and 0 otherwise. we set $K=p/10$. The description and results for three other structures we consider (tridiagonal, hubs and cliques negative) and trace plots to assess MCMC mixing are deferred to Supplementary Section S.3.
@@
For each $(n,p)$ pair, we draw $50$ replicate data sets from $\mathcal{N}_{p}(0,\boldsymbol{\Theta}_{0}^{-1})$. Each MCMC run consists of $10,000$ iterations with the first $5,000$ discarded as burn-in and uses a single 2.0 GHz AMD EPYC Rome CPU core with 16 GB of memory, and a maximum run time limit of 4 hours. Our de novo implementation of the RT sampler is in Rcpp with underlying C++ code, whereas the existing R implementations of the cyclical sampler for GHS (Srivastava and Bhadra,, 2018) and GHSL (Sagar et al., 2024a, ) are used. Given this difference in implementation, we take particular care to clarify the basis of our runtime comparison in the subsequent parts. The standard half-Cauchy prior is used on the global parameter $\tau$ for both GHS and GHSL.
@@
For each replication, we record the following posterior summaries of $\boldsymbol{\Theta}$: (i) $\|\widehat{\boldsymbol{\Theta}}\|_{F}$ under both samplers, (ii) $\|\widehat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}_{0}\|_{F}$ under both samplers, and (iii) $\|\widehat{\boldsymbol{\Theta}}_{\mathrm{RT}}-\widehat{\boldsymbol{\Theta}}_{\mathrm{Cyclical}}\|_{F}$, where $\|\cdot\|_{F}$ denotes the Frobenius norm. Tables 1 and 2 report means, variances, and the 25%, 50%, 75% empirical quartiles averaged over the $50$ replications, with Monte Carlo standard errors in parentheses, for the GHS and GHSL priors respectively. For both priors, the Frobenius norms of the estimated precision matrices obtained from the two samplers are numerically close across all $(n,p)$ settings. Similarly, the deviations from the truth show negligible differences between the two samplers. Moreover, the pairwise deviations between RT and the baseline sampler remain small for all $(n,p)$, indicating that both samplers produce highly consistent posterior estimates. Given that the RT and cyclical samplers target the same posterior under two different reparameterizations, these results are expected, and indeed a statistically significant difference would indicate a problem with either or both the samplers.
@@
Table 1: Estimation results for the RT and cyclical samplers under the cliques positive structure for the GHS prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (“–”) indicates the cyclical sampler exceeded the 44-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 28.42(1.28) 28.20(1.23) 20.71(1.38) 20.54(1.32) 1.95(0.50)
Var 3.10(0.36) 3.01(0.36) 10.51(0.09) 10.55(0.09) 0.80(0.19)
25% 27.79(1.29) 27.62(1.24) 20.38(1.38) 20.26(1.32) 2.01(0.65)
50% 28.24(1.27) 28.03(1.23) 20.55(1.38) 20.38(1.33) 1.99(0.63)
75% 29.30(1.28) 29.02(1.24) 21.48(1.38) 21.25(1.33) 2.31(0.54)
(127, 200) Mean 39.91(1.24) 39.39(1.23) 28.89(1.32) 28.49(1.32) 2.88(0.63)
Var 3.64(0.37) 3.70(0.42) 14.74(0.08) 14.87(0.10) 1.15(0.29)
25% 39.01(1.25) 38.61(1.24) 28.39(1.32) 28.10(1.32) 2.94(0.70)
50% 39.66(1.23) 39.17(1.22) 28.66(1.32) 28.30(1.31) 3.02(0.79)
75% 41.06(1.25) 40.46(1.24) 29.88(1.34) 29.41(1.33) 3.47(0.81)
(137, 300) Mean 48.63(1.33) - 35.08(1.42) - -
Var 4.00(0.32) - 18.00(0.06) - -
25% 47.54(1.32) - 34.46(1.41) - -
50% 48.34(1.32) - 34.82(1.42) - -
75% 49.97(1.34) - 36.23(1.44) - -
Table 2: Estimation results for the RT and cyclical samplers under the cliques positive structure for the GHSL prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (“–”) indicates the cyclical sampler exceeded the 44-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 28.11(1.31) 28.15(1.24) 20.46(1.40) 20.49(1.33) 2.12(0.54)
Var 3.20(0.39) 3.04(0.35) 10.54(0.11) 10.56(0.09) 0.94(0.22)
25% 27.50(1.32) 27.57(1.24) 20.18(1.41) 20.21(1.33) 2.26(0.70)
50% 27.94(1.31) 27.98(1.23) 20.33(1.40) 20.33(1.33) 2.34(0.73)
75% 28.98(1.31) 28.97(1.24) 21.24(1.40) 21.20(1.33) 2.55(0.66)
(127, 200) Mean 39.67(1.27) 39.44(1.23) 28.73(1.35) 28.51(1.31) 2.59(0.57)
Var 3.74(0.41) 3.62(0.40) 14.75(0.09) 14.84(0.09) 1.14(0.28)
25% 38.78(1.28) 38.65(1.24) 28.23(1.36) 28.11(1.32) 2.64(0.68)
50% 39.43(1.27) 39.22(1.22) 28.52(1.35) 28.31(1.31) 2.83(0.76)
75% 40.83(1.29) 40.51(1.24) 29.74(1.37) 29.42(1.33) 3.16(0.72)
(137, 300) Mean 48.21(1.23) - 34.73(1.33) - -
Var 4.05(0.32) - 17.98(0.07) - -
25% 47.11(1.24) - 34.10(1.34) - -
50% 47.93(1.23) - 34.48(1.33) - -
75% 49.58(1.23) - 35.91(1.34) - -
@@
Figure 1: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHS prior under the cliques positive structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,150,200,250,300\}$, with $n\approx 24\log(p)$. At $p=300$, the cyclical sampler exceeded the 4-hour wall time limit, so its raw runtime is capped at 4 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
@@
Figure 2: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHSL prior under the cliques positive structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,150,200,250,300\}$, with $n\approx 24\log(p)$. At $p=300$, the cyclical sampler exceeded the 4-hour wall time limit, so its raw runtime is capped at 4 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
@@
Figures 1 and 2 compare the raw and relative runtimes for the RT and cyclical samplers under the GHS and GHSL priors, respectively. It is worth explaining why we present both raw and scaled runtime plots in each figure, where the latter is relative to the runtime at $p=100$, computed for each method against its own implementation. Although an end user is primarily concerned with the raw runtime, this quantity cannot be easily compared across different implementations to assess algorithmic scaling. An algorithm that scales poorly in the dimension $p$, but implemented in C++, could still outperform a better scaling algorithm, but implemented in R, well into fairly large dimensions in terms of raw runtime. Hence, for a fair comparison of the algorithms, we also present the scaled runtime as a reasonable proxy that is approximately independent of the implementation. As a consequence, all scaled runtime plots start from a $y$-axis value of one at $p=100$. The figures demonstrate that we comfortably outperform Li et al., (2019)’s and Sagar et al., 2024a ’s algorithms in both raw and scaled runtimes, but it is only in the light of the scaled runtime that we claim algorithmic superiority. The relative advantage of RT increases with $p$: at $p=100$, the RT sampler is approximately $5.1$ times faster than Cyclical in raw runtime, whereas at $p=300$ it is at least $7.7$ times faster in raw runtime and $1.5$ times faster in scaled runtime. This is consistent with the $\mathcal{O}(p^{3})$ and $\mathcal{O}(p^{4})$ scalings of the RT and cyclical samplers when $n=o(p^{1/2})$. To summarize, taking the results of Li et al., (2019) and Sagar et al., 2024a as ground truths, the tables serve to demonstrate the correctness of our reparameterized MCMC, while the figures demonstrate its improved scalability. Computer code implementing the proposed approach and example simulations are freely available on Github at: https://github.com/gao702/RTsampler.
@@ 5 Analysis of TCGA Breast Cancer Data
We analyze the Cancer Genome Atlas (TCGA,, 2013) breast cancer data set (release: 6 May 2017), preprocessed following Zhang et al., (2019). Considering the genes only in the KEGG breast cancer pathway yields $p=139$ features. We focus on the basal-like subtype with RNA-sequencing measurements, yielding $n=90$ samples. All variables are standardized prior to analysis. We draw $5,000$ posterior samples, discarding the first $1,000$ as burn-in, under both GHS and GHSL priors using cyclical and RT schemes.
@@
Table 3: Runtime (seconds) comparing RT vs. cyclical samplers under GHS and GHSL priors for the breast cancer RNA sequencing data ($p=139,n=90$). MCMC run uses 5,0005{,}000 iterations with the first 1,0001{,}000 discarded as burn-in.
@@
6 Conclusions and Future Works
The objective of the current work is to present a reverse telescoping sampling scheme under element-wise graphical priors, of the form of Equations (1)–(2). Although for demonstration purposes we choose two such priors, the graphical horseshoe and the graphical horseshoe-like, the proposed sampler is generically applicable to all element-wise priors of the above form. In the high-dimensional regime $n=o(p^{1/2})$, the proposed approach attains a per iteration complexity of $\mathcal{O}(p^{3})$, providing a one order-of-magnitude improvement over the $\mathcal{O}(p^{4})$ cyclical sampler. Considerable numerical results are presented to support our claim of improved algorithmic scaling.

### v2

@@ 4 Numerical Experiments
We compare the proposed reverse telescoping (RT) MCMC sampler with the cyclical MCMC sampler under the graphical horseshoe (GHS, Li et al.,, 2019) and graphical horseshoe-like (GHSL, Sagar et al., 2024a, ) priors in terms of estimation results and computational efficiency. To ensure $n=o(p^{1/2})$, we set $n=24\log p$ (rounded to the nearest integer) for $p=100,200,300,400,600,800$; resulting in the following problem dimensions: $(n,p)\in\{(110,100),(127,200),(137,300),(143,400),(153,600),(160,800)\}$. Recall the theory suggests in this regime, the RT sampler achieves $\mathcal{O}(\max\{n^{2}p^{2},p^{3}\})=\mathcal{O}(p^{3})$ per iteration complexity, whereas the complexity of the cyclical sampler remains $\mathcal{O}(p^{4})$. Our choice of not introducing a new prior, and instead working with the existing GHS and GHSL priors is intentional, as this allows us to isolate the computational performance of the proposed RT and the existing cyclical samplers, without confounding this with statistical performance due to different model choices. Nevertheless, the proposed sampler is applicable to all priors of the form of Equations (1)–(2).
@@
The true precision matrix $\boldsymbol{\Theta}_{0}$ is set as in Li et al., (2019) and Sagar et al., 2024b . The diagonal elements are set to one and the off-diagonal elements are set to follow the hubs structure, which means the $p$ rows and columns are partitioned into $K$ disjoint groups $\{\mathbf{G}_{k}\}_{k=1}^{K}$. The off-diagonal entry is set to $0.25$ if $i\neq j$ and $i,j\in\mathbf{G}_{k}$, and 0 otherwise. We set $K=p/10$, with all groups having equal size.
@@
For each $(n,p)$ pair, we draw $50$ replicate data sets from $\mathcal{N}_{p}(0,\boldsymbol{\Theta}_{0}^{-1})$. Each MCMC run consists of $10,000$ iterations with the first $5,000$ discarded as burn-in and uses a single 2.0 GHz AMD EPYC Rome CPU core with 32 GB of memory, and a maximum runtime limit of 12 hours. Our de novo implementation of the RT sampler is in Rcpp with underlying C++ code, whereas the existing R implementations of the cyclical sampler for GHS (Srivastava and Bhadra,, 2018) and GHSL (Sagar et al., 2024a, ) are used. Given this difference in implementation, we take particular care to clarify the basis of our runtime comparison in the subsequent parts. The standard half-Cauchy prior is used on the global parameter $\tau$ for both GHS and GHSL.
4.1 Estimation and Timing Results
@@
For each replication, we record the following posterior summaries of $\boldsymbol{\Theta}$: (i) $\|\widehat{\boldsymbol{\Theta}}\|_{F}$ under both samplers, (ii) $\|\widehat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}_{0}\|_{F}$ under both samplers, and (iii) $\|\widehat{\boldsymbol{\Theta}}_{\mathrm{RT}}-\widehat{\boldsymbol{\Theta}}_{\mathrm{Cyclical}}\|_{F}$, where $\|\cdot\|_{F}$ denotes the Frobenius norm. Tables 1 and 2 report means, variances, and the 25th, 50th, 75th empirical percentiles averaged over the $50$ replications, with Monte Carlo standard errors in parentheses, for the GHS and GHSL priors respectively. For both priors, the Frobenius norms of the estimated precision matrices obtained from the two samplers are numerically close across all $(n,p)$ settings. Similarly, the deviations from the truth show negligible differences between the two samplers. Moreover, the pairwise deviations between RT and the baseline sampler remain small for all $(n,p)$, indicating that both samplers produce highly similar posterior estimates. Given that the RT and cyclical samplers target the same posterior under two different reparameterizations, these results are expected, and indeed a statistically significant difference would indicate a problem with either or both the samplers.
@@
Table 1: Estimation results for the RT and Cyclical samplers under the hubs structure for the GHS prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on up to $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (–) indicates the cyclical sampler exceeded the $12$-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 15.00(0.50) 14.36(0.39) 5.86(0.50) 5.20(0.39) 1.08(0.20)
Var 0.60(0.03) 0.53(0.02) 10.10(0.02) 10.14(0.01) 0.15(0.02)
25% 13.40(0.45) 12.85(0.35) 5.00(0.39) 4.52(0.28) 1.24(0.21)
50% 14.87(0.49) 14.25(0.39) 5.77(0.48) 5.17(0.37) 1.09(0.21)
75% 16.60(0.55) 15.87(0.42) 7.20(0.57) 6.43(0.44) 1.23(0.23)
(137, 300) Mean 26.55(0.58) 24.96(0.43) 10.53(0.60) 8.95(0.43) 2.29(0.24)
Var 0.91(0.03) 0.76(0.01) 17.61(0.02) 17.69(0.01) 0.29(0.03)
25% 23.99(0.53) 22.59(0.39) 8.96(0.48) 7.74(0.31) 2.57(0.23)
50% 26.36(0.57) 24.82(0.42) 10.40(0.58) 8.94(0.41) 2.32(0.24)
75% 29.10(0.61) 27.34(0.45) 12.71(0.66) 10.91(0.49) 2.69(0.28)
(160, 800) Mean 44.25(0.54) - 17.90(0.53) - -
Var 1.39(0.03) - 28.85(0.02) - -
25% 40.26(0.50) - 15.32(0.42) - -
50% 43.97(0.54) - 17.70(0.51) - -
75% 48.22(0.58) - 21.35(0.59) - -
Table 2: Estimation results for the RT and Cyclical samplers under the hubs structure for the GHSL prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on up to $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (–) indicates the cyclical sampler exceeded the $12$-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 14.82(0.51) 14.33(0.39) 5.70(0.50) 5.18(0.39) 0.96(0.21)
Var 0.62(0.04) 0.53(0.02) 10.08(0.03) 10.14(0.01) 0.16(0.03)
25% 13.19(0.46) 12.83(0.35) 4.85(0.37) 4.51(0.28) 1.11(0.20)
50% 14.68(0.50) 14.23(0.39) 5.63(0.47) 5.16(0.38) 0.96(0.22)
75% 16.44(0.57) 15.85(0.43) 7.06(0.58) 6.41(0.45) 1.11(0.26)
(137, 300) Mean 26.32(0.96) 25.10(0.48) 10.33(0.90) 9.07(0.51) 2.16(0.69)
Var 0.97(0.08) 0.76(0.02) 17.56(0.05) 17.69(0.01) 0.34(0.07)
25% 23.65(0.86) 22.72(0.45) 8.69(0.69) 7.81(0.39) 2.37(0.58)
50% 26.11(0.93) 24.97(0.48) 10.20(0.82) 9.05(0.49) 2.20(0.62)
75% 28.94(1.04) 27.50(0.50) 12.59(1.00) 11.06(0.55) 2.63(0.73)
(160, 800) Mean 41.72(1.71) - 15.78(1.46) - -
Var 1.30(0.12) - 28.90(0.07) - -
25% 37.82(1.55) - 13.46(1.13) - -
50% 41.49(1.66) - 15.75(1.33) - -
75% 45.60(1.85) - 19.12(1.63) - -
@@
Figure 1: Average raw runtimes in minutes (left panel) and runtimes relative to $p=100$ (right panel), for the GHS prior under the hubs structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,200,300,400,600,800\}$, with $n=24\log p$. At $p>400$, the cyclical sampler exceeded the 12-hour runtime limit. The RT sampler is implemented in Rcpp and the cyclical sampler in R.
@@
Figure 2: Average raw runtimes in minutes (left panel) and runtimes relative to $p=100$ (right panel), for the GHSL prior under the hubs structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,200,300,400,600,800\}$, with $n=24\log p$. At $p>400$, the cyclical sampler exceeded the 12-hour runtime limit. The RT sampler is implemented in Rcpp and the cyclical sampler in R.
@@
Figures 1 and 2 compare the raw and relative runtimes for the RT and cyclical samplers under the GHS and GHSL priors, respectively. It is worth explaining why we present both raw and scaled runtime plots in each figure, where the latter is relative to the runtime at $p=100$, computed for each method against its own implementation. Although an end user is primarily concerned with the raw runtime, this quantity cannot be easily compared across different implementations to assess algorithmic scaling. An algorithm that scales poorly in the dimension $p$, but implemented in C++, could still outperform a better scaling algorithm, but implemented in R, well into fairly large dimensions in terms of raw runtime. Hence, for a fair comparison of the algorithms, we also present the scaled runtime as a reasonable proxy that is approximately independent of the implementation. As a consequence, all scaled runtime plots start from a $y$-axis value of one at $p=100$. The figures demonstrate that we comfortably outperform Li et al., (2019)’s and Sagar et al., 2024a ’s algorithms in both raw and scaled runtimes, but it is only in the light of the scaled runtime that we claim algorithmic superiority. The relative advantage of RT increases with $p$: at $p=100$, the RT sampler is approximately $5.1$ times faster than the cyclical sampler in raw runtime, whereas at $p=400$ it is $8.2$ times faster in raw runtime. Similar patterns hold for the scaled runtimes. This is consistent with the $\mathcal{O}(p^{3})$ and $\mathcal{O}(p^{4})$ scalings of the RT and cyclical samplers when $n=o(p^{1/2})$. To summarize, taking the results of Li et al., (2019) and Sagar et al., 2024a as ground truths, the tables serve to demonstrate the correctness of our reparameterized MCMC, while the figures demonstrate its improved scalability.
4.2 Assessment of MCMC Mixing
We assess mixing via effective sample size per unit time (ESS/min), computed from the scalar log-likelihood trace across MCMC iterations. As $p$ increases, ESS/min decreases for both methods, which is expected. However, at a given $p$, the RT sampler consistently demonstrates 2–5 times higher ESS/min, when results are available for both methods.
Table 3: Mean (sd) of ESS per minute computed from the log-likelihood trace under the hubs structure for the GHS and GHSL priors, based on 50 replicates. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in and a thinning interval of 5. A dash (–) indicates the cyclical sampler exceeded the $12$-hour runtime limit and no results were obtained.
GHS GHSL
Method $p=100$ $p=300$ $p=800$ $p=100$ $p=300$ $p=800$
Cyclical 10.44 (4.20) 0.10 (0.04) 9.54 (3.06) 0.09 (0.04)
RT 48.48 (16.50) 0.57 (0.30) 0.38 (0.79) 22.32 (6.24) 0.28 (0.16) 0.03 (0.01)
4.3 Additional Results and Code Availability
Results for three other structures for the data-generating $\boldsymbol{\Theta}_{0}$ we consider (tridiagonal, cliques negative and cliques positive) and trace plots to assess MCMC mixing are deferred to Supplementary Section S.3. The additional results are consistent with the above findings, that is, the RT sampler produces samples of comparable quality but achieves an order of magnitude speed-up over the cyclical sampler, under all these alternate structures. Computer code implementing the proposed approach in R package RTsampler and example simulations are freely available on GitHub at: https://github.com/gao702/RTsampler.
@@ 5 Analysis of TCGA Breast Cancer Data
We analyze the Cancer Genome Atlas (TCGA,, 2013) breast cancer data set (release: 6 May 2017), preprocessed following Zhang et al., (2019). Considering the genes only in the KEGG breast cancer pathway yields $p=139$ features. We focus on the basal-like subtype with RNA sequencing measurements, yielding $n=90$ samples. All variables are standardized prior to analysis. We draw $5,000$ posterior samples, discarding the first $1,000$ as burn-in, under both GHS and GHSL priors using cyclical and RT schemes.
@@
Table 4: Runtime (seconds) comparing RT versus cyclical samplers under GHS and GHSL priors for the breast cancer RNA sequencing data ($p=139,n=90$). MCMC run uses 5,0005{,}000 iterations with the first 1,0001{,}000 discarded as burn-in.
@@
6 Conclusions and Future Work
The objective of the current work is to present a reverse telescoping sampling scheme under element-wise graphical priors, of the form of Equations (1)–(2). Although for demonstration purposes we choose two such priors, the graphical horseshoe and the graphical horseshoe-like, the proposed sampler is generically applicable to all element-wise priors of the above form. In the high-dimensional regime $n=o(p^{1/2})$, the proposed approach attains a per iteration complexity of $\mathcal{O}(p^{3})$, providing a one order of magnitude improvement over the $\mathcal{O}(p^{4})$ cyclical sampler. Considerable numerical results are presented to support our claim of improved algorithmic scaling.